In class activity

On June 1, 2020, Jill Bow and Aisha Adams formed a partnership to open a gluten-free commercial bakery, contributing $281,000 cash and $362,000 of equipment, respectively. The partnership also assumed responsibility for a $41,000 note payable associated with the equipment. The partners agreed to share profits as follows: Bow is to receive an annual salary allowance of $151,000, both are to receive an annual interest allowance of 5% of their original capital investments, and any remaining profit or loss is to be shared 40/60 (to Bow and Adams, respectively). On November 20, 2020, Adams withdrew cash of $101,000. At year-end, May 31, 2021, the Income Summary account had a credit balance of $390,000. On June 1, 2021, Peter Williams invested $121,000 and was admitted to the partnership for a 20% interest in equity.
Answer #1

Part 1

No.

Date

Account titles and explanation

Debit

Credit

a

June 1, 2020

Cash

281000

Equipment

362000

J. Bow, Capital

281000

A. Adams, Capital

321000

Notes Payable

41000

To record formation of partnership

b

November 20, 2020

A. Adams, Withdrawals

101000

Cash

101000

To record withdrawal by partner.

c

May 31, 2021

Income summary

390000

J. Bow, Capital

206830

A. Adams, Capital

183170

To record closing of net income to capital

d

June 1, 2021

Cash

121000

J. Bow, Capital (202400-121000)*20%

16280

A. Adams, Capital (202400-121000)*80%

65120

P. Williams, Capital (1012000*20%)

202400

To record the admission of new partner

J. Bow

A. Adams

Total

Net income

390000

Salary allowance:

Bow

151000

Interest allowances:

Bow (5% on $281000)

14050

Adams (5% on $321000)

16050

Total salaries and interest allocation

165050

16050

181100

Balance of income to be allocated

208900

Balance allocated 40/60:

Bow (20% × $208900)

41780

Adams (80% × $208900)

167120

Total allocated 40/60

(208900)

Balance of income

0

Shares of the partners

206830

183170

390000

Total equity prior to admission of new partner = (281000+206830)+(321000-101000+183170) = 891000

Total equity after admission of new partner = 891000+121000= 1012000

Part 2

J. Bow, Capital (281000+206830-16280)

$471550

A. Adams, Capital (321000-101000+183170-65120)

$338050

P. Williams, Capital

$133500